a capacitor has 4.33 x 10^-7 C of charge on it when 3.45 V is applied. how much energy is stored in the capacitor
Answer #1

here,

the capacitance , C = 4.33 * 10^-7 C

the potential difference , V = 3.45 V

the energy stored in the capacitor , U = 0.5 * C * V^2

U = 0.5 * 4.33 * 10^-7 * 3.45^2 J

U = 2.58 * 10^-6 J